OpenStudy (anonymous):

Verify: Csc^2xtan^2x-1=tan^2x

OpenStudy (raden):

notice that : csc^2 x = 1/sin^2 x tan^2 x = sin^2 x/cos^2 x you can start by using the identities above

OpenStudy (raden):

so, the left side can be Csc^2xtan^2x-1 = 1/sin^2 x * sin^2 x/cos^2 x cancel the sin^2 x we get Csc^2xtan^2x-1 = 1/cos^2 x and remember the identities bellow : 1/cos^2 x = sec^2 x and sec^2 - 1 = tan^2 x therefore, Csc^2xtan^2x-1 = sec^2 x - 1 = tan^2 x = RHS

OpenStudy (anonymous):

Ah okay thank you very much! (:

OpenStudy (raden):

wlcm ;)